1. Solve for the price of a forward contract on a generic

asset that expires on September 10 whose spot price

as of June 10 is $45, assuming that the annually

compounded risk-free rate is 6.01 percent.
Answer #1

Ans : From June 10 to September 10 = Time period = T = 3/12

S0 = $ 45

F(0,T) = S0(1+r)T

= 45 *(1+6.01%)3/12

= 45 *(1.0601)1/4

= 45 * 1.01469

= $ 45.66139